Here is the data collected from my projectile motion and conservation of energy lab on an even plane. The Theoretical range that I calculated was R=1.68. I am to GRAPH Range vs. Initial launch angle and then CALCULATE the % difference between the measured range(?) and the theoretical range (1.68) for the initial launch angle of 45 degrees. PLEASE show work

Table 3: Even y Plane ; = 45 AX1 AX2H AXSH
Answer #1
\Theta=450
\Delta x_1 106.5
\Delta x_2 107.0
\Delta x_3 107.5
\Delta x_4 108.0
\Delta x_5 108.0
\Delta x_{avg} 107.4

\frac{\Delta x_{avg}}{\Theta =45^{0}}=\frac{107.4}{45}=2.386

\frac{\Delta x_{th}}{\Theta =45^{0}}=1.68

Percentage.Relative . Error =\frac{2.386-1.68}{1.68}*100=42.1%